OpenStudy (anonymous):

what is the simplified form of (2/x^2+x) - (1/x) ?? a.(x-1)/(x(x+1) b.(1-x)/(x(x+1) c.(3-x)/(x(x-1) d.(x+2)/(x(x-1) Thank you in advance !

OpenStudy (anonymous):

a/b - c/d = (ad - bc)/bd Here a = 2, b = x^2+x, c = 1, d = x so the common denominator b times d is (x(x+1)). What is ad - bc?

OpenStudy (anonymous):

sorry denominator = x(x^2 + x)

OpenStudy (anonymous):

I'm sorry but I don't know what ad-bc would be.

OpenStudy (anonymous):

If we define a = 2, b = x^2+x, c = 1, d = x then your first equation is the same as a/b - c/d OK?

OpenStudy (anonymous):

yes I'm following

OpenStudy (anonymous):

So then ad = 2x and bc = .. and ad - bc = ..

OpenStudy (anonymous):

x^2+x*1= x^2+x right since its times 1?

OpenStudy (anonymous):

So now do the subtraction

OpenStudy (anonymous):

Well, it's 2x - (x^2 + x) = 2x - x^2 - x = x - x^2 = x(1-x). That's your numerator

OpenStudy (anonymous):

so would the final answer be (1-x)/(x(x+1)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

We had to do some factoring at the end.

OpenStudy (anonymous):

Thank you very much!